if f ‘(x) = 8(x)^3 – 2x , f(2) = 8 find f(x)

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

f'(x) = 8x^3-2x
\frac{df(x)}{dx} = 8x^3-2x
df(x) = (8x^3-2x)dx
\int df(x) = \int (8x^3-2x)dx
f(x) = 2x^4 - x^2 + c
f(2) = 2.2^4 - 2^2 + c
8 = 32 - 4 + c
c = -20
f(x) = 2x^4 - x^2 - 20